Mathematics 10 — Exercise Review · Q1 (10)
Full question
P(E)=0.07. Probability of not E? (a)0.95 (b)0.89 (c)0.93 (d)0.90
Step-by-step solution
1−0.07=0.93.
Answer
(c)0.93.
P(E)=0.07. Probability of not E? (a)0.95 (b)0.89 (c)0.93 (d)0.90
1−0.07=0.93.
(c)0.93.